# Inverse sampling intensity weighting for preferential sampling adjustment

### v1

@@ 2.1.3 Vecchia approximation
The Vecchia approximation is a specific case of CL based on the observation that the joint distribution can be decomposed as the product of conditional distributions. Let $\textbf{Y}=(Y_{1},...,Y_{n})^{\top}\in\mathbb{R}^{n}$ and $p:\{1,...,n\}\to\{1,...,n\}$ be a permutation mapping which reorders Y. We define the history of variable $Y_{j}$ as a random subvector $\textbf{Y}_{h(j)}$ where $h(j)=\{l\in\{1,...,n\}:p(l)<p(j)\}$. The joint density of Y can then be refactored as
@@
$\begin{split}\log\mathcal{L}_{V}(\boldsymbol{\psi};\textbf{y})&=w_{1}\log f(y_{p(1)};\boldsymbol{\psi})\\ &+\sum_{i=2}^{n}w_{1i}f(y_{p(i)},\textbf{y}_{q(i)};\boldsymbol{\psi})\\ &-\sum_{i=2}^{n}w_{2i}\log f(\textbf{y}_{q(i)};\boldsymbol{\psi}).\end{split}$ (5)
@@
While simple in concept, the Vecchia approximation requires careful selection of three key hyperparameters: 1) the size of the conditioning set, 2) which variables to include in each conditioning set, and 3) the ordering of the variables as determined by $p$. We denote the maximum size of any $q(i)$ as $m$. In all later sections, we choose $m=20$ based on a clear case of diminishing returns in inference and prediction for $m>20$ observed empirically in Datta et al., (2016). Default settings in two implementations of the Vecchia approximation (GPvecchia and GpGp R packages) are $m=20$ and $m=30$, respectively (Katzfuss and Guinness,, 2021; Guinness,, 2021). For the choice of which variables to include in the neighborhood $q(i)$, we follow the recommendation given in Vecchia, (1988) by choosing the nearest neighbors to $y_{p(i)}$ measured by Euclidean distance. Ordering for the Vecchia approximation requires additional specification in multidimensional settings due to the lack of natural ordering in observations of a spatial point process. We use the maxmin ordering scheme, which picks a location $p(i)$ sequentially by maximizing the distance to the nearest point in $\{y_{p(1)},y_{p(2)},...,y_{p(i-1)}\}$ and has been shown to approximate the true distribution better than other ordering methods (Guinness,, 2018; Katzfuss and Guinness,, 2021).
Theoretical justification of the Vecchia approximation in spatial statistics is relatively light compared to that of the MLE and CL. Typical approaches to proving fixed domain asymptotics fail due to the loss of stationarity in the parent process introduced by the Vecchia approximation (Zhang et al.,, 2024). However, it is impossible to deny its excellent empirical performance in applications (Heaton et al.,, 2019). A key advantage the Vecchia approximation possesses over typical CL is that (4) and (5) correspond to a valid joint probability distribution for Y. As a result, the Kullback-Leibler (KL) divergence of the Vecchia approximation with respect to the true distribution can be computed and has been shown to be a nonincreasing function of $m$ when the conditioning sets $q(i)$ are chosen as the nearest neighbors (Katzfuss et al.,, 2020; Huser et al.,, 2023).
The Vecchia approximation achieves a good balance of computational complexity and efficiency. It clearly outperforms other CL methods in terms of what order correlations are being accounted for, scalability, and probabilistic validity. This is the main reason for considering its use in ISIW methods compared to CL, a point we expand on in later sections.

### v2

@@ 2.2 Vecchia approximation
The Vecchia approximation is a specific case of composite likelihood (varin_overview_2011) based on the observation that the joint distribution of a random vector can be decomposed as the product of conditional distributions. Let $\textbf{Y}=(Y_{1},...,Y_{n})^{\top}\in\mathbb{R}^{n}$ and $p:\{1,...,n\}\to\{1,...,n\}$ be a permutation mapping which reorders Y. Here $f(\cdot;\boldsymbol{\psi})$ denotes the probability density of the corresponding subvector of $\mathbf{Y}$ under parameter vector $\boldsymbol{\psi}$. We define the history of variable $Y_{j}$ as a random subvector $\textbf{Y}_{h(j)}$ where $h(j)=\{l\in\{1,...,n\}:p(l)<p(j)\}$. The joint density of Y evaluated at a fixed value $\mathbf{y}$ can then be refactored as
@@
$\begin{split}\log\mathcal{L}_{V}(\boldsymbol{\psi};\textbf{y})&=w_{1}\log f(y_{p(1)};\boldsymbol{\psi})\\ &+\sum_{i=2}^{n}w_{1i}\log f(y_{p(i)},\textbf{y}_{q(i)};\boldsymbol{\psi})\\ &-\sum_{i=2}^{n}w_{2i}\log f(\textbf{y}_{q(i)};\boldsymbol{\psi}).\end{split}$ (3)
@@
While simple in concept, the Vecchia approximation requires careful selection of three key hyperparameters: 1) the size of the conditioning set, 2) which variables to include in each conditioning set, and 3) the ordering of the variables as determined by $p$. We denote the maximum size of any $q(i)$ as $m$. In all later sections, we choose $m=20$ based on a clear case of diminishing returns in inference and prediction for $m>20$ observed empirically in datta_hierarchical_2016. Default settings in two implementations of the Vecchia approximation (GPvecchia and GpGp R packages) are $m=20$ and $m=30$, respectively (katzfuss_general_2021; guinness_gaussian_2021). For the choice of which variables to include in the neighborhood $q(i)$, we follow the recommendation given in vecchia_estimation_1988 by choosing the nearest neighbors to $y_{p(i)}$ measured by Euclidean distance. Ordering for the Vecchia approximation requires additional specification in 2D settings due to the lack of natural ordering in observations of a multidimensional point process. We use the maxmin ordering scheme, which picks a location $p(i)$ sequentially by maximizing the distance to the nearest point in $\{y_{p(1)},y_{p(2)},...,y_{p(i-1)}\}$ and has been shown to approximate the true distribution better than other ordering methods (guinness_permutation_2018; katzfuss_general_2021).
The Vecchia approximation strikes a favorable balance between computational efficiency and statistical accuracy. Compared with other composite likelihood approaches, it accommodates higher-order dependence, scales well to large datasets, and corresponds to a valid joint density. This is the main reason for considering its use in ISIW methods compared to univariate and pairwise composite likelihoods from previous studies.
2.3 Kriging
For many spatial analyses, the main goal is to predict values at unobserved locations. A standard spatial prediction tool is kriging, or Gaussian process regression, a function approximation technique with point and variance estimate defined as
$\operatorname{Var}\big\{\hat{S}(\textbf{X}_{0})\big\}$ $=C_{\theta}(\textbf{X}_{0},\textbf{X}_{0})-C_{\theta}(\textbf{X}_{0},\textbf{X}_{n})^{\top}\boldsymbol{\Sigma}_{n}^{-1}(\theta)C_{\theta}(\textbf{X}_{0},\textbf{X}_{n}),$
where $\textbf{X}_{0}$ represents unobserved locations to be predicted. In practice, since the parameters are unknown, all parameters in (4) are replaced with estimates obtained from any of the aforementioned estimation methods.
It can be shown under the assumption of NPS that kriging is the best linear unbiased predictor (BLUP) (cressie_statistics_2015). However, this optimality no longer holds under preferential sampling, where even the true values of $\boldsymbol{\psi}$ used in the kriging equations yield suboptimal predictions (dinsdale_methods_2019). This is because the optimal predictor is the conditional expectation $\mathbb{E}\big\{S(\mathbf{X}_{0})\mid\mathbf{Y},\mathbf{X}\big\}$, whose distribution is no longer Gaussian as in the fixed case. Nevertheless, kriging can still be used to generate predictions under these non-standard conditions. One potential approach is to solve for the kriging weights under the model defined in (1). However this optimization does not admit a closed-form expression. We therefore turn to ISIW as a practical approach for finding GP mean and covariance parameters that yield an approximation to the optimal linear kriging predictor under PS, even though these parameters may not coincide with the true latent random field.
3 Inverse sampling intensity weighting (ISIW)
We now describe the ISIW procedure for PS adjustment. ISIW is a two-stage approach, wherein we first estimate the sampling intensity at each of the observation locations X. In the second stage, we input the (estimated) inverse sampling intensities as weights into a weighted likelihood adjustment for a chosen composite likelihood. For spatial interpolation, the resulting adjusted parameter estimates are then substituted in the kriging equations in (4).
3.1 Sampling intensity estimation
The essence of ISIW is to account for the dependence between the response and observation locations by using the vector of estimated sampling intensities instead of the full likelihood of the observation process. The heuristic motivation for ISIW is to reweight observations by inverse sampling intensity to approximate a representative, non-preferential spatial sample of the study region.
Methods to estimate a spatially varying intensity of a point process can be divided into parametric and nonparametric approaches. Domain knowledge can inform the parametric form of the intensity either through choice of model or covariates, but oftentimes this information is unavailable and nonparametric estimation is required. The nonparametric approaches follow the kernel smoothing approach discussed in diggle_kernel_1985 which we will refer to as kernel intensity estimators (KIE). Let $k$ be a $d$-dimensional kernel function from $\mathbb{R}^{d}\to\mathbb{R}^{+}$, which is a symmetric probability density function. Given a bandwidth size $h>0$ and edge correction factor $w_{h}(\cdot,\cdot)$, the KIE estimate at a point $\mathbf{x}$ is given by
$\hat{\lambda}(\textbf{x};h)=h^{-d}\sum_{\textbf{s}\in\textbf{X}\cap\mathcal{D}}k\left(\frac{\lVert\textbf{x}-\textbf{s}\rVert}{h}\right)w_{h}(\textbf{x},\textbf{s})^{-1},\textbf{x}\in\mathcal{D}.$
The key hyperparameter for KIE is the bandwidth size $h$. Bandwidth selection is a well-studied problem, with methods ranging from high bandwidth, smooth intensity estimators to low bandwidth, flexible intensity estimators. Each has its place in the bias-variance tradeoff, with higher bandwidths exhibiting higher bias with lower variance and lower bandwidths vice versa. We experiment with several bandwidth selection strategies implemented in the spatstat R package (Table S1).
While KIEs provide valid estimates of the first-order intensity based solely on a realization of the point pattern, a potential limitation for PS adjustment is that they do not incorporate information from the observed responses $\mathbf{Y}$. Under the SLP model in (1), for example, the sampling intensity depends on the latent process, which also drives the observed response values. Incorporating both response values and spatial locations may therefore yield improved estimates of the sampling intensity relative to approaches based only on the locations.
To explore this idea, we study an extension of KIE that incorporates the response through a spatial covariate using the kernel-based approach of baddeley_nonparametric_2012. We also considered a random forest–based intensity estimation method proposed by biscio_nonparametric_2025; however, its results were very similar to those obtained with KIE so we do not present them here. Specifically, we use a predicted version of $\mathbf{Y}$ as an additional covariate in the intensity estimation. Because KIEs with spatial covariates require covariate values to be available at all grid locations, we obtain this covariate by kriging using parameters estimated via MLE. Although this predicted surface is biased, as has been extensively discussed, PS induces systematic discrepancies between the true process and the estimated surface. If such structure is present, the nonparametric relationship between the covariate and the sampling intensity may still be learnable. We refer to this approach as “ISIW KIE COV” and the approach using just the locations as “ISIW KIE.”
As with other inverse weighting procedures, a key concern for ISIW is the presence of extreme weights. Observation locations in sparsely sampled regions, which occur more frequently under PS designs, can receive very large weights in the resulting likelihood and may lead to numerical instability during optimization. Two commonly used approaches from the inverse probability weighting literature are trimming and winsorization. Trimming removes observations deemed to be outliers, whereas winsorization preserves the sample size by capping extreme values at a specified threshold.
We adopt winsorization to mitigate the impact of extreme weights, as trimming observations may discard too many informative locations, thereby weakening the effectiveness of the PS adjustment. Our winsorization procedure proceeds as follows. First, we estimate the sampling intensities and take their inverse. An upper quantile is then selected as the winsorization threshold, and all values exceeding this threshold are set equal to it. Finally, the weights are normalized to sum to the total number of observations $n$. Denote the $\tau$-level empirical quantile of the estimated inverse sampling intensities $\hat{\lambda}^{-1}$ as $W_{\tau}$. The resulting estimated weight for the $i$th observation is
$\hat{w}_{i}=n\cdot\frac{\min\{\hat{\lambda}(\mathbf{x}_{i})^{-1},W_{\tau}\}}{\sum_{\mathbf{s}\in\mathbf{X}}\min\{\hat{\lambda}(\mathbf{s})^{-1},W_{\tau}\}}.$ (5)
3.2 Weighted likelihood and estimation
ISIW applies weights proportional to the inverse sampling intensity to each event in a likelihood factored as a product of densities. While such weighting is straightforward in settings with independent observations, the appropriate weighting scheme is less clear in the presence of spatial dependence and correlation among observations. Existing applications of ISIW in the PS literature include univariate marginal (vedensky_look_2023) and pairwise difference CLs (schliep_correcting_2023). These approaches yield intuitive weighting schemes by enforcing independence across subsets of events in the joint density. For reference, the weighted univariate marginal CL follows
$\begin{split}\log\mathcal{L}_{WM}(\mu;\textbf{y})=\sum_{i=1}^{n}w_{i}\log f(y_{i};\mu),\end{split}$
while the pairwise difference CL follows
$\begin{split}\log\mathcal{L}_{WD}(\theta;\textbf{y})=\sum_{i=1}^{n}\sum_{j=i+1}^{n}w_{i}w_{j}\log f(y_{i}-y_{j};\theta).\end{split}$
However, the univariate marginal CL cannot estimate covariance parameters $\theta$ while the pairwise difference CL cannot estimate $\mu$ directly, limiting its impact on PS adjustment. In addition, both approaches ignore higher-order spatial dependence, which is essential for accurately approximating Gaussian process models.
To address these gaps, we explore the application of ISIW to the Vecchia approximation. Let the weights be defined as in (5). We initially considered weighting the likelihood as
to maintain the probabilistic interpretation of the inverse weighting for the component conditional densities. However, numerical issues arise due to taking the product of several weights, greatly increasing the chance of extreme values. As a more stable approach, we approximate the true weight using
$\begin{split}\log\mathcal{L}_{WV}(\boldsymbol{\psi};\textbf{y})&=w_{p(1)}\log f(y_{p(1)};\boldsymbol{\psi})\\ &+\sum_{i=2}^{n}w_{p(i)}\log f(y_{p(i)},\textbf{y}_{q(i)};\boldsymbol{\psi})\\ &-\sum_{i=2}^{n}w_{p(i)}\log f(\textbf{y}_{q(i)};\boldsymbol{\psi}).\end{split}$ (6)
We also considered the pairwise marginal composite likelihood as a candidate for ISIW, as it naturally accommodates weighting while allowing estimation of $\mu$. However, preliminary analyses indicated that ISIW applied to the Vecchia approximation outperformed the pairwise marginal CL across all evaluation metrics (Table S2). We therefore focus the remainder of the paper on ISIW combined with Vecchia.
The ISIW parameters can be numerically estimated by standard optimization procedures. We used the L-BFGS-B routine as implemented in the optim package in the R language. In our experience, other derivative-free optimization procedures such as the Nelder-Mead algorithm also work well. Initial values for parameters to be estimated were selected with general rules of thumb from the GPVecchia R package and all other optimization parameters were set to their default settings.